OpenStudy (amarab):

simplify the expression? would like help, not an answer. -6+i/-5+i

OpenStudy (asnaseer):

do you know what the conjugate of a complex number \((a+ib)\) is?

OpenStudy (amarab):

a-bi?

OpenStudy (asnaseer):

correct. now notice what happens if you multiple a complex number by its conjugate:\[(a+bi)(a-bi)=a^2-(bi)^2=a^2+b^2\]

OpenStudy (asnaseer):

i.e. the imaginary parts of a complex number disappear when you multiply it by its conjugate

OpenStudy (amarab):

okay thanks! just what i needed

OpenStudy (asnaseer):

so the idea here is to multiply your fraction with something that would result in the denominator losing its imaginary part.

OpenStudy (asnaseer):

ok - glad you got it! :)

OpenStudy (amarab):

okay so i did... -6+i(-5-i) over -5+i(-5-i) then got 30+6i+-5i+-i^2 or 30+2i^2 but to the bottom... i ended up with 25+2i^2 and i don't know what to do after that??

OpenStudy (amarab):

it looks like 30+2i^2/25+2i^2 but i know it's wrong

OpenStudy (anonymous):

i^2=-1

OpenStudy (asnaseer):

\[\frac{-6+i}{-5+i}=\frac{-6+i}{-5+i}\times\frac{-5-i}{-5-i}=\frac{(-6+i)(-5-i)}{(-5+i)(-5-i)}\]\[\qquad=\frac{30+6i-5i-i^2}{25+5i-5i-i^2}=\frac{30+i-i^2}{25-i^2}\]Then (as suggested by @irene22988) use the fact that \(i^2=-1\) to simplify further